Pittman Company is a small but growing manufacturer of telecommunications equipment. The company has no sales force of its owI say its just plain robbery, retorted Karl. “And I also say its time we dumped those guys and got our own sales force. C2. Assume that Pittman Company decides to continue selling through agents and pays the 20% commission rate. Determine the dolRequired 1 Required 2 Required 3 Required 4 Assume that Pittman Company decides to continue selling through agents and pays tComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Determine theRequired i Required 2 Required 3 Required 4 Compute the degree of operating leverage that the company would expect to have at
Answer #1
Solution 1:
15% commission 20% commission own sales force
Sales 20000000 20000000 20000000
Variable expenses:
Manufacturing 9000000 9000000 9000000
Commission 3000000 4000000 1500000
Total variable expenses 12000000 13000000 10500000
Contribution margin 8000000 7000000 9500000
Variable expenses ratio 60.00% 65.00% 52.50%
Contribution margin ratio 40.000% 35.000% 47.500%
Fixed expenses:
Manufacturing overhead 2800000 2800000 2800000
Marketing 140000 140000 3140000
Administrative 1960000 1960000 1868000
Interest 700000 700000 700000
Total fixed expenses 5600000 5600000 8508000
Income before income taxes 2400000 1400000 992000
Income taxes(30%) 720000 420000 297600
Net income 1680000 980000 694400
Break Even Point in dollars (Total Fixed Expenses/Controbution margin ratio) 14000000 16000000 17911579
Solution 2:
Income before tax as per income statement 2400000
Total Fixed Expenses (when 20% commission) 5600000
Total 8000000
/ Contribution margin ratio (When 20% commission) 35.000%
Sales required to generate target profit 22857143
Solution 3:
Let Sales Volume be "X"
Total Cost when commission 20% = Total Cost for Own sales Force
0.6500*X +5600000 = 0.5250*X +8508000
0.125*X = 2908000
X = 2908000/0.125
X = 23264000
Hence, Sales volume required is $23,264,000.
Solution 4:
15% commission 20% commission own sales force
Contribution margin 8000000 7000000 9500000
Income before taxes 2400000 1400000 992000
Degree of Operating Leverage (Contribution margin/Income before taxes) 3.33 5.00 9.58